OpenStudy (anonymous):

find derivative \[\frac{ \log_{x} 2 }{ \sqrt{x} } + 2^{-x}\]

OpenStudy (anonymous):

\[\frac{ \log_{x} 2 }{ \sqrt{x} } + 2^{-x}\]

OpenStudy (raden):

use division rule to find derivative of log_x (2)/sqrt(x)

OpenStudy (anonymous):

can you help?

OpenStudy (raden):

which parts did u stuck

OpenStudy (anonymous):

log_x (2)/sqrt(x)

OpenStudy (raden):

use the formula y=u/v ---> y'=(u'v-uv')/v^2

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

what about the last part

OpenStudy (raden):

deivative of y=a^x is y'=a^x * ln(a)

OpenStudy (raden):

a is constant